OpenStudy (anonymous):

I am stuck on a problem...I wish I knew the name of it, but I will type it out as best as I can. It's finding common function, and I have to plot its points. The problem is 9 to the exponent of 6 squareroot x / 7 (so it's a fraction with the 7 as a denominator)

OpenStudy (anonymous):



OpenStudy (anonymous):

I don't know even where to start on this...

OpenStudy (anonymous):

I need to find two points

OpenStudy (anonymous):

pick numbers that have sixth roots to make your life easy x = 0 x = 1 x = 64

OpenStudy (anonymous):

i have no idea what 3^6 is so stick with those few

OpenStudy (anonymous):

what does sixth roots mean?

OpenStudy (anonymous):

ok, never mind. I understand that far...but I don't understand how I would work the rest of the problem to come up with my points

OpenStudy (anonymous):

i picked 0^6 = 0 1^6 = 1 2^6 = 64 if you only need two points that should be enough

OpenStudy (anonymous):

What do I do with the 9 and the 7? That's what throwing me off the most.

OpenStudy (anonymous):

0 gives you 0 so you can ignore them x = 1 you get \[\frac{9}{7}\] x = 64 you get \[\frac{18}{7}\]

OpenStudy (anonymous):

\[\frac{9\sqrt[6]{64}}{7}=\frac{9\times 2}{7}=\frac{18}{7}\]

OpenStudy (anonymous):

oooh!! perfect, thank you! That explains what I needed to know

OpenStudy (anonymous):

yw